Evaluate csc (48 degrees 48 minutes)?
Answer #1

have to find csc( 48 degrees 48 minute)

now, 48 minute = 48/60 degree ( since 1 minute = 1/60 degree)

= 0.8 degree

therefore 48 degree 48 minute = 48.8 degree

now, csc( 48 degree 48 minute) = csc (48.8 degree)

= 1/sin(48.8 degree)

= 1/0.752

= 1.33 ( approximately)